OpenStudy (anonymous):

2x4(4x2 + 3x + 1) (4x – 3)(2x2 – 7x + 1) (x2 + 4x – 3)(2x2 + x + 6)

OpenStudy (anonymous):

Simplify and show work please

OpenStudy (anonymous):

This is a very easy question but its very tedious essentially, all you have to do is use foil, where you multiply each number in the second bracket by every number in the first bracket So, it'd go like this: x^2(2x^2) +x^2(x) + X^2(-6) =2x^4 + x^3 +6x^2) And you'd continue this pattern, getting: 4x(2x^2) + 4x(x) -3(6)= (8x^3 + 4x^2 + 24x) and (-6x^2 - 3x -18) for -3 After you get these three parts, you simply collect like terms, (2x^4 + x^3 +6x^2) + (8x^3 + 4x^2 + 24x) + (-6x^2 - 3x -18) =2x^4 + 9x^3 +4x^2 +21x - 18 and the ^'s signify exponents